OpenStudy (anonymous):

Fun time: Simple Puzzle: Marry's Father has 10 daughters: Their names are: 1. Nanerir210 2. Nanerir211 3. Nanerir213 4. Nanerir214 5. Nanerir217 6. Nanerir218 7. Nanerir222 8. Nanerir223 9. Nanerir228 Name of his 10th daughter will be : ??

OpenStudy (anonymous):

i think Nanerir238

OpenStudy (anonymous):

Well Done.. Any Logic ??

OpenStudy (anonymous):

Yeah @sauravshakya is right..

OpenStudy (anonymous):

i think the first daughter and 4 th daughter and seventh daughter are at gapes of increasing multiples of 4

OpenStudy (anonymous):

Nice one......

OpenStudy (anonymous):

no it should be Nanerir229 :P

OpenStudy (anonymous):

Nice logic @Mitul

OpenStudy (anonymous):

@sami-21 you can take more chances..

OpenStudy (anonymous):

ty @waterineyes

OpenStudy (anonymous):

@mathslover any idea ??

mathslover (mathslover):

The middle row is the main fault of my idea .. 1. Nanerir210 2. Nanerir211 3. Nanerir213 4. Nanerir214 5. Nanerir217 6. Nanerir218 7. Nanerir222 8. Nanerir223 9. Nanerir228 1)+4=4) 2)+6=5) 3)+5=6) 4)+8=7) 5)+6=8) 9)+10=228 But yes it will not lead to have the correct answer...

mathslover (mathslover):

please wait .. lemme think about this

OpenStudy (anonymous):

Take your time.. You have all the time..

OpenStudy (vishweshshrimali5):

nice question

OpenStudy (anonymous):

marry

OpenStudy (anonymous):

Am I missing something? Diffs are just 1-2-1-3-1-4-1-5 Next diff is 1: 9. Nanerir229

OpenStudy (anonymous):

I mean 10: Nanerir 229

OpenStudy (anonymous):

You don't deserve a medal.. Ha ha ha...

OpenStudy (lgbasallote):

lol. that's harsh

OpenStudy (anonymous):

@lgbasallote looking nice in this picture..